Solve by rearranging the equations first. Then use the multiplication principle and then use the elimination method:
3x=8y+11 x+ 6y - 8 = 0
This is confusing.

Answer by mananth(16946)

3x=8y+11
shift the terms to one side and the constants to the other side
-8y
3x-8y=11......................1
x+ 6y - 8 = 0
+8
x+6y=8........................2
..
multiply equation (2) by -3
-3x-18y=-24
..
add this equation to (1)
-3x+3x-18y-8y=-24+11
-26y=-13
/-26
y= -13/-26
y = 1/2
...
plug value of y in equation (2)
x+6*0.5 = 8
x+3=8
-3
x=8-3
x=5
..
m.ananth@hotmail.ca